If 1000 g of Sodium will reacts with 100 ppm of Oxygen, How much Sodium Oxide will form?
4 Na + O2 = 2 Na2O
1
Expert's answer
2020-02-27T09:01:35-0500

n(Na)=m/M(Na)=1000/23=43.48 mol

n(Na2O)=1/2n(Na)=43.48/2=21.74 mol

m(Na2O)=n*M(Na2O)=21.74*62=1347.8 g

Answer: 1347.8 g.